OpenStudy (anonymous):

Logx + Log(x- 2) = Log3x

OpenStudy (shamim):

Do u know log(a)+log(b)=log(ab)

OpenStudy (anonymous):

yes

OpenStudy (shamim):

So we can write log(x)+log(x-2)=log(x(x-2)) Right?

OpenStudy (shamim):

I think u know If log(a)=log(b)

OpenStudy (shamim):

Then we can write a=b

OpenStudy (anonymous):

ok

OpenStudy (shamim):

log(x(x-2))=log(3x)

OpenStudy (shamim):

x=?

OpenStudy (anonymous):

so wouldn't it be x^2 - 2x?

OpenStudy (shamim):

Yes

OpenStudy (shamim):

U r correct!!

OpenStudy (anonymous):

so just solve for x...

OpenStudy (shamim):

Ya

OpenStudy (shamim):

x^2-2x=3x Right?

OpenStudy (shamim):

x=?

OpenStudy (anonymous):

x^2 = 5x

OpenStudy (shamim):

Correct!!

OpenStudy (anonymous):

sqr 5x

OpenStudy (shamim):

No

OpenStudy (shamim):

x^2-5x=0 Right?

OpenStudy (anonymous):

you add 5x to both sides right

OpenStudy (shamim):

x^2=5x x^2-5x=5x-5x x^2-5x=0 Right?

OpenStudy (anonymous):

yes thats right ....